OpenStudy (anonymous):

Four times the sum of a number and 15 is at least 120. Let x represent the number. Find all possible values for x ???? HELP!

OpenStudy (zzr0ck3r):

4(x + 15)=< 120 then solve

OpenStudy (zzr0ck3r):

make since?

OpenStudy (anonymous):

I don't know equations yet

OpenStudy (zzr0ck3r):

I dont think there is any other way to do this

OpenStudy (zzr0ck3r):

what are you doing in class that they asked you this question?

OpenStudy (anonymous):

idk that was the only thing they wrote besides solve lol

OpenStudy (anonymous):

im doing online work

OpenStudy (zzr0ck3r):

hmm well to do this you need to work with an equation

OpenStudy (zzr0ck3r):

4 times the sum of a number and 15 is at least 120 4 * (x+15) >= 120

OpenStudy (zzr0ck3r):

my first reply should say 4*(x+15) >= 120 where >= meens greater than or equal to

OpenStudy (anonymous):

is it x3 15?

OpenStudy (zzr0ck3r):

4*(x+15)>=120 4x+60>=120 4x>= 60 x >= 15

OpenStudy (zzr0ck3r):

I dont know if your 3 is a > but dont use it:0